# SCBench: A KV Cache-Centric Analysis of Long-Context Methods

### v1

@@ 1 Introduction
Despite the benefit, Long-context inputs also present unique challenges for LLM inference due to high computational costs and memory demands. This has led to the development of efficient long-context solutions leveraging sparsity in various stage of KV cache. In this paper, we introduce an unified analysis framework for efficient long-context methods in a KV cache centric perspective, consisting of the four essential stages of KV cache: 1) KV cache generation, 2) KV cache compression, 3) KV cache retrieval, and 4) KV cache loading. First, KV cache generation, aka prefilling, processes the input prompt and produce the KV cache to be used in decoding. In this stage, sparse attention methods are proposed to reduce the complexity of the attention operation (Child et al., 2019; Beltagy et al., 2020; Jiang et al., 2024). Second, KV cache compression techniques prune KV states to reduce the memory costs in decoding (Xiao et al., 2024b; Li et al., 2024c). Third, KV cache retrieval aims to skip KV cache generation of an incoming request, and instead retrieves and reuses KV cache from history KV cache pool for more efficient inference (Zheng et al., 2024; Yao et al., 2024a). At last, KV cache loading aims to load only partial of the KV cache for each decoding step to save the memory and computing cost (Tang et al., 2024; Liu et al., 2024b).
However, these methods are only evaluated on single-request benchmarks (Hsieh et al., 2024; Zhang et al., 2024a; Kamradt, 2023; Li et al., 2024a), which fail to covers the full lifecycle of KV cache in real applications. Typically, real-world applications often require reusing prompt memory (i.e., KV cache) and involving multiple requests or multi-round interactions (Qin et al., 2024). The reuse of KV cache, known as prefix caching, is already a crucial component in popular inference frameworks (Zheng et al., 2024; vLLM, 2024) and used by LLM providers (Gemini, 2024; Claude, 2024; OpenAI, 2024b; Azure, 2024). In addition, testing with multiple requests is especially crucial for the long-context methods mentioned earlier, as many achieve efficiency through query-conditioned compression. For instance, Arora et al. (2024) reports that Mamba’s compression of previous information based on the current query can prevent it from answering follow-up queries.
@@
With SCBench, we conduct an extensive KV cache-centric analysis in the four stages as shown in Fig. 1 (details in Sec. 2). Specifically, we evaluate 13 long-context methods across four stages and eight categories on eight open-source long-context LLMs, including Llama-3.1-8B/70B (Dubey et al., 2024), Qwen2.5-72B/32B (Team, 2024), Llama-3-8B-262K (Gradient, 2024), GLM-4-9B-1M (GLM et al., 2024), Codestal Mamba (team, 2024), and Jamba-1.5-mini (Lieber et al., 2024). These methods span gated linear RNNs (e.g., Codestal Mamba), hybrid models (e.g., Jamba-1.5), sparse attention (e.g., A-shape, Tri-shape, MInference (Jiang et al., 2024)), prompt compression (e.g., LLMLingua-2 (Pan et al., 2024)), KV cache dropping (e.g., StreamingLLM (Xiao et al., 2024b), SnapKV (Li et al., 2024c)), KV cache quantitation (e.g., KIVI (Liu et al., 2024e)), semantic retrieval (e.g., CacheBlend (Yao et al., 2024a)), and KV cache loading (e.g., Quest (Tang et al., 2024), RetrievalAttention (Liu et al., 2024b)), as detailed in Table 1. Additionally, we introduce Tri-shape, a novel, training-free sparse attention method that demonstrates improved first-turn performance in our evaluations.
Our experimental results reveal the following insights: 1) Sub-$O(n)$ memory is almost infeasible in multi-turn decoding, as shown in Fig. 3. Sparse decoding methods (sub-$O(n)$ memory) perform well on the first query but lose accuracy in subsequent requests. In contrast, sparse encoding methods ($O(n)$ memory with $O(n^{2})$computation during pre-filling) can approximate full attention accuracy across multiple queries. 2) Task performance shows varying decline trends, as illustrated in Fig. 3b. Sparse KV cache methods excel in tasks requiring global information, whereas $O(n)$ memory is essential for tasks involving exact match retrieval. 3) All long-context methods experience performance degradation as the budget decreases, as shown in Fig. 4. However, sub-$O(n)$ memory methods exhibit a significant performance drop at a 1/4 compression rate. Methods such as RetrievalAttention and KIVI, which maintain $O(n)$ memory with sparse decoding, sustain higher performance even at higher compression rates. 4) Long-generation scenarios exhibit distribution shift issues, as generation length and the number of rounds increase, the importance distribution of the KV cache changes significantly. This out-of-distribution (OOD) issue leads to performance degradation, even for $O(n)$ memory methods like RetrievalAttention, as observed in extended tasks, as shown in Fig. 3.
@@ 2 A KV Cache-Centric Perspective on Long-Context Methods
A-shape (Xiao et al., 2024b) Sparse Attention $O(n)$ $O(kn)$ $O(nm)$
@@
StreamingLLM (Xiao et al., 2024b) KV Cache Dropping $O(k)$ $O(n^{2})$ $O(km)$
@@
Recently, a series of works (Gu & Dao, 2024; Xiao et al., 2024b; Jiang et al., 2024) have explored various strategies to reduce the inference cost of long-context LLMs, enabling their application (Jimenez et al., 2024; OpenAI, 2024a) to downstream tasks at a lower computational expense. In long-context LLM inference, the KV cache plays a pivotal role by effectively reducing the computational overhead during the decoding phase. This importance has led to the development of numerous system-level optimizations (Sheng et al., 2023; Qin et al., 2024) focused on KV cache management and scheduling.
@@
    KV Cache Generation: This stage optimizes the efficient generation of KV cache during inference. Techniques include sparse attention (e.g., A-shape, Tri-shape, MInference (Jiang et al., 2024)), SSM or hybrid approaches (e.g., Mamba (Gu & Dao, 2024; Dao & Gu, 2024; Lieber et al., 2024)), and prompt compression (e.g., LLMLingua-2 (Pan et al., 2024)).
@@
    KV Cache Compression: After generation, the KV cache is compressed before being stored. Methods include KV cache dropping (e.g., StreamingLLM (Xiao et al., 2024b), SnapKV (Li et al., 2024c)) and KV cache quantization (e.g., KIVI (Liu et al., 2024e)).
@@
    KV Cache Loading: This stage dynamically loads the KV cache and computes sparse attention, from KV cache storage (e.g., VRAM, DRAM, SSD, or RDMA) to GPU on-chip SRAM, including Quest (Tang et al., 2024), RetrievalAttention (Liu et al., 2024b), and MagicPIG (Chen et al., 2024).
@@
We also introduce a novel training-free sparse attention method, Tri-shape, with improved first-turn accuracy, as shown in Fig. 5. Specifically, in addition to retaining the sink token and local window regions preserved by A-shape, Tri-shape also retains the last window query region, forming a triangular pattern for sparse attention during the pre-filling stage. The motivation arises from the observation from our SCBench that A-shape with dense decoding exhibits significant performance improvement after multiple requests. Tri-shape can notably enhance performance in both turn-0 and multi-request scenarios, as detailed in Sec. 4. Furthermore, it preserves the ability of LLMs to follow instructions, as demonstrated in the case study in §F. This is an example of how our SCBench can provide insights for novel algorithm innovation. Notably, some recent concurrent works (Acharya et al., 2024) have also proposed similar patterns to accelerate the long-context pre-filling stage.
@@ 3 Benchmark Building
SCBench comprises 12 tasks covering four long-context abilities: string retrieval, semantic retrieval, global information processing, and multi-tasking, across two shared context modes—multi-turn and multi-request. These tasks span various domains, including code, retrieval, question answering, summarization, in-context learning, multi-hop tracing, and multi-tasking, as shown in Fig. 2b. In total, SCBench includes 931 multi-turn sessions with 4,853 queries, averaging 5 turns per session. Task statistics are provided in Table 2, with examples and configurations in Table 3. Below, we detail the construction of our benchmark.
@@ 3.1 Long-context Task Details
The most fundamental requirement for long-context LLMs is the ability to identify and retrieve information relevant to a specific query from a lengthy, potentially noisy input. To evaluate this, string retrieval tasks are widely used, where models must retrieve a specific string based on given conditions (Hsieh et al., 2024; Zhang et al., 2024a). Our benchmark incorporates complexity analysis, similar to approaches used in algorithmic problem-solving, such as LeetCode, to design three distinct tasks with varying levels of difficulty. Additionally, by varying the position of the target string, our benchmark further evaluates how well models utilize the full extent of their claimed context window (Kamradt, 2023).
(i) Retrieve.KV: Given a large JSON object containing numerous key-value pairs, the models must accurately retrieve the value corresponding to a specified key (Liu et al., 2024d). The random KVs in this task present significant challenges for long-context LLMs, as the input is often incompressible, requiring strict $O(n)$ space to store. This makes it particularly useful for testing the fuzziness of memory in long-context methods, especially in KV Cache usage. In each session, five KV pairs are retrieved, with the target KVs evenly distributed across the full length of the input.
(ii) Retrieve.Prefix-Suffix: Given a large list of variable-length strings, the models must accurately retrieve a string with a specific prefix and suffix. This task is particularly challenging, as the models need to implement complex functions to match both the prefix and suffix (similar to a prefix tree, with a computational cost of $O(\sum{w_{i}}^{2})$, where $w_{i}$ represents the length of the $i$-th string [^1] https://leetcode.com/problems/prefix-and-suffix-search/). The presence of distractors that share either the prefix or suffix, but not both, prevents models from relying on simple lookup mechanisms or induction heads (Olsson et al., 2022) to solve the task effectively.
(iii) Retrieve.MultiHop: This task, first proposed in RULER (Hsieh et al., 2024), is designed to evaluate the multi-hop tracing capabilities of LLMs within a long input prompt. It requires models to capture and memorize changes in key information from the input context, making it ideal for testing long-context methods in KV cache reuse. Five multi-hop variable assignment chains are embedded throughout the context, and each turn in the test session requires the models to retrieve the exact multi-hop chain, i.e., all variables assigned to a specific value.
@@
In addition to string retrieval, many real-world long-context applications require semantic understanding beyond simple string matching, such as retrieving a function based on textual descriptions or answering questions from a long document. These tasks are crucial in SCBench, as lossy long-context methods may struggle to abstract or comprehend information in multi-request scenarios.
(i) Code.RepoQA: This task requires the model to retrieve a specific function (including the function name, input parameters, and full implementation) from a long chunk of source code based on a precise natural language description. Unlike the original RepoQA benchmark (Liu et al., 2024c), our inputs are extended to 64K tokens, with target functions evenly selected based on their position in the codebase. The function descriptions were generated using GPT-4 based on the functions themselves. Additionally, we expanded the range of repositories and programming languages in our test to include Python, C++, Java, PHP, Rust, Go, and TypeScript, compared to the original RepoQA. Each test session involves a GitHub repository, and the model is required to retrieve one function per turn, with a total of 5 turns per session.
(ii) En.QA, Zh.QA, En.MultiChoice: These three tasks are extended from InfiniteBench (Zhang et al., 2024a), which provides high-quality, human-annotated QA tests based on fictional novels to eliminate the influence of external knowledge. These tasks require models to locate and process information within lengthy inputs, performing reasoning through aggregation or filtering to derive answers. There are two primary types of questions: 1) Aggregation involves compiling scattered information throughout the input. An example question is, “How much money in total did A spend on food?” 2) Filtering equires identifying specific information from a larger set. An example question is, “What color dress did A wear when A met B for the second time?” In SCBench, we combine QA pairs that share the same input context to create shared context test sessions.
@@
In addition to retrieval, some long-context tasks require leveraging and aggregating global context information, such as summarization, statistical tasks, and in-context learning (Yu et al., 2020; Srivastava et al., 2023; Hao et al., 2022). Our benchmark includes three relevant tasks to assess how well different long-context methods handle global information in multi-request settings.
(i) Many-shot ICL: We use datasets from Big-Bench Hard (Srivastava et al., 2023) to evaluate many-shot in-context learning (ICL) capabilities. This includes three sub-tasks: date understanding, salient error translation detection, and tracking seven shuffled objects. We construct many-shot ICL contexts shared across different turns within a test session. All three sub-tasks are presented as multiple-choice questions with four options provided.
(ii) Math.Find: We extended the math find task from InfiniteBench (Zhang et al., 2024a), expanding from finding only maximum value to multiple statistical values. Given a large array, LLMs are required to find the minimum or median values. LLMs must effectively comprehend global long-context information, perform comparisons, and carry out statistical operations to answer the questions.
(iii) En.Sum: This task uses concatenated academic papers from arXiv as input, with document lengths ranging from 8K to 20K tokens. Ground truth summaries were generated using GPT-4, which was prompted to produce concise one-sentence summaries for each document. The average length of the ground truth summaries is 654 tokens. The target documents for each turn are evenly distributed across the full context length.
@@
In real-world applications, LLMs often handle multiple tasks within a single session using a shared input context. For instance, users might request both summarization and content retrieval simultaneously. To reflect this, we include two multi-tasking tasks in SCBench:
(i) Mix.Sum+NIAH: This task combines document summarization with the Needle in a Haystack (Kamradt, 2023) task using a shared input prompt. A random "needle" is evenly inserted into the En.Sum task’s input (concatenated academic papers). The model alternates between summarization and NIAH retrieval in each test session.
(ii) Mix.RepoQA+KV: This task combines the RepoQA task with KV retrieval using a shared input prompt. Multiple KV pairs are evenly inserted into the RepoQA input (a long chunk of source code). A total of 100 KV pairs are included, with four target KVs and the rest as distractors. The model alternates between RepoQA and KV retrieval in each test session.

### v2

@@ 1 Introduction
Despite the benefit, Long-context inputs also present unique challenges for LLM inference due to high computational costs and memory demands. This has led to the development of efficient long-context solutions leveraging sparsity in various stage of KV cache. In this paper, we introduce an unified analysis framework for efficient long-context methods in a KV cache centric perspective, consisting of the four essential stages of KV cache: 1) KV cache generation, 2) KV cache compression, 3) KV cache retrieval, and 4) KV cache loading. First, KV cache generation, aka prefilling, processes the input prompt and produce the KV cache to be used in decoding. In this stage, sparse attention methods are proposed to reduce the complexity of the attention operation (Child et al., 2019; Beltagy et al., 2020; Jiang et al., 2024). Second, KV cache compression techniques prune KV states to reduce the memory costs in decoding (Xiao et al., 2024; Li et al., 2024c). Third, KV cache retrieval aims to skip KV cache generation of an incoming request, and instead retrieves and reuses KV cache from history KV cache pool for more efficient inference (Zheng et al., 2024; Yao et al., 2024a). At last, KV cache loading aims to load only partial of the KV cache for each decoding step to save the memory and computing cost (Tang et al., 2024; Liu et al., 2024b).
However, these methods are only evaluated on single-request benchmarks (Hsieh et al., 2024; Zhang et al., 2024a; Kamradt, 2023; Li et al., 2024a), which fail to covers the full lifecycle of KV cache in real applications. Typically, real-world applications often require reusing prompt memory (i.e., KV cache) and involving multiple requests or multi-round interactions (Qin et al., 2025). The reuse of KV cache, known as prefix caching, is already a crucial component in popular inference frameworks (Zheng et al., 2024; vLLM, 2024) and used by LLM providers (Gemini, 2024; Claude, 2024; OpenAI, 2024b; Azure, 2024). In addition, testing with multiple requests is especially crucial for the long-context methods mentioned earlier, as many achieve efficiency through query-conditioned compression. For instance, Arora et al. (2024) reports that Mamba’s compression of previous information based on the current query can prevent it from answering follow-up queries.
@@
With SCBench, we conduct an extensive KV cache-centric analysis in the four stages as shown in Fig. 1 (details in §2). Specifically, we evaluate 13 long-context methods across four stages and eight categories on eight open-source long-context LLMs, including Llama-3.1-8B/70B (Dubey et al., 2024), Qwen2.5-72B/32B (Team, 2024), Llama-3-8B-262K (Gradient, 2024), GLM-4-9B-1M (GLM et al., 2024), Codestal Mamba (team, 2024), and Jamba-1.5-mini (Lieber et al., 2024). These methods span gated linear RNNs (e.g., Codestal Mamba), hybrid models (e.g., Jamba-1.5), sparse attention (e.g., A-shape, Tri-shape, MInference (Jiang et al., 2024)), prompt compression (e.g., LLMLingua-2 (Pan et al., 2024)), KV cache dropping (e.g., StreamingLLM (Xiao et al., 2024), SnapKV (Li et al., 2024c)), KV cache quantitation (e.g., KIVI (Liu et al., 2024e)), semantic retrieval (e.g., CacheBlend (Yao et al., 2024a)), and KV cache loading (e.g., Quest (Tang et al., 2024), RetrievalAttention (Liu et al., 2024b)), as detailed in Table 1. Additionally, we introduce Tri-shape, a novel, training-free sparse attention method that demonstrates improved first-turn performance in our evaluations.
Our experimental results reveal the following insights: 1) Sub-$O(n)$ memory is almost infeasible in multi-turn decoding, as shown in Fig. 3. Sparse decoding methods (sub-$O(n)$ memory) perform well on the first query but lose accuracy in subsequent requests. In contrast, sparse encoding methods ($O(n)$ memory with $O(n^{2})$ computation during pre-filling) can approximate full attention accuracy across multiple queries. 2) Task performance declines at varying rates, as illustrated in Fig. 3b. Sparse KV cache methods excel in tasks requiring global information, while $O(n)$ memory is crucial for exact match retrieval tasks. 3) All long-context methods degrade as the budget decreases, as shown in Fig. 4. However, sub-$O(n)$ memory methods experience a sharp performance drop at a 1/4 compression rate. Methods like RetrievalAttention and KIVI, which maintain $O(n)$ memory with sparse decoding, sustain higher performance even under higher compression rates. 4) Long-generation scenarios exhibit distribution shift issues. As generation length and the number of rounds increase, the KV cache’s importance distribution changes significantly. This out-of-distribution (OOD) issue leads to performance degradation, even for $O(n)$ memory methods like RetrievalAttention, as shown in Fig. 3.
@@ 2 A KV Cache-Centric Perspective on Long-Context Methods
A-shape (Xiao et al., 2024) Sparse Attention $O(n)$ $O(kn)$ $O(nm)$
@@
StreamingLLM (Xiao et al., 2024) KV Cache Dropping $O(k)$ $O(n^{2})$ $O(km)$
@@
Recently, a series of works (Gu & Dao, 2024; Xiao et al., 2024; Jiang et al., 2024) have explored various strategies to reduce the inference cost of long-context LLMs, enabling their application (Jimenez et al., 2024; OpenAI, 2024a) to downstream tasks at a lower computational expense. In long-context LLM inference, the KV cache plays a pivotal role by effectively reducing the computational overhead during the decoding phase. This importance has led to the development of numerous system-level optimizations (Sheng et al., 2023; Qin et al., 2025) focused on KV cache management and scheduling.
@@
    KV Cache Generation: This stage optimizes the efficient generation of KV cache during inference. Techniques include sparse attention (e.g., A-shape, Tri-shape, MInference (Jiang et al., 2024), NSA (Yuan et al., 2025), MoBA (Lu et al., 2025)), SSM or hybrid approaches (e.g., Mamba (Gu & Dao, 2024; Dao & Gu, 2024; Lieber et al., 2024)), and prompt compression (e.g., LLMLingua-2 (Pan et al., 2024)).
@@
    KV Cache Compression: After generation, the KV cache is compressed before being stored. Methods include KV cache dropping (e.g., StreamingLLM (Xiao et al., 2024), SnapKV (Li et al., 2024c)) and KV cache quantization (e.g., KIVI (Liu et al., 2024e)).
@@
    KV Cache Loading: This stage dynamically loads the KV cache and computes sparse attention, from KV cache storage (e.g., VRAM, DRAM, SSD, or RDMA) to GPU on-chip SRAM, including Quest (Tang et al., 2024), RetrievalAttention (Liu et al., 2024b), and MagicPIG (Chen et al., 2025).
@@
We introduce Tri-shape, a novel training-free sparse attention method that improves first-turn accuracy (Fig. 5). Unlike A-shape, which retains only the sink token and local window, Tri-shape also preserves the last window query region, forming a triangular sparse attention pattern for pre-filling. Motivated by our SCBench findings that A-shape with dense decoding improves after multiple requests, Tri-shape enhances both turn-0 and multi-request performance (§4) while maintaining LLM instruction-following ability (§G). Notably, recent concurrent work (Acharya et al., 2024) has explored similar patterns for accelerating long-context pre-filling.
@@ 3 Benchmark Building
SCBench features 12 tasks assessing four long-context abilities: string retrieval, semantic retrieval, global information processing, and multi-tasking, across two shared context modes—multi-turn and multi-request. These tasks span diverse domains, including code, retrieval, question answering, summarization, in-context learning, and multi-hop tracing (Fig. 2b). In total, SCBench includes 931 multi-turn sessions with 4,853 queries, averaging 5 turns per session. Task statistics are provided in Table 2, with examples and configurations in Table 3. Below, we describe the benchmark construction.
@@ 3.1 Long-context Task Details
The core requirement for long-context LLMs is retrieving relevant information from lengthy, potentially noisy inputs. String retrieval tasks are commonly used for this evaluation (Hsieh et al., 2024; Zhang et al., 2024a). Our benchmark applies complexity analysis, inspired by algorithmic problem-solving (e.g., LeetCode), to design three tasks with varying difficulty levels. Additionally, by shifting the target string’s position, we assess how well models utilize their full context window (Kamradt, 2023).
(i) Retrieve.KV: Given a large JSON object with numerous key-value pairs, models must accurately retrieve the value for a specified key (Liu et al., 2024d). The random KVs pose challenges for long-context LLMs, as the input is often incompressible, demanding strict $O(n)$ space for storage. This task is particularly useful for evaluating memory fuzziness in long-context methods. Each session retrieves five KV pairs, with target KVs evenly distributed across the input.
(ii) Retrieve.Prefix-Suffix: Given a large list of variable-length strings, models must retrieve a string matching a specified prefix and suffix. This task is particularly challenging, requiring complex functions akin to a prefix tree, with a computational cost of $O(\sum{w_{i}}^{2})$, where $w_{i}$ is the length of the $i$-th string [^1] https://leetcode.com/problems/prefix-and-suffix-search/. The presence of distractors sharing only the prefix or suffix prevents models from relying on simple lookups or induction heads (Olsson et al., 2022) for effective retrieval.
(iii) Retrieve.MultiHop: First introduced in RULER (Hsieh et al., 2024), this task evaluates LLMs’ multi-hop tracing capabilities within long input prompts. Models must track and recall key information changes, making it ideal for testing long-context methods in KV cache reuse. Five multi-hop variable assignment chains are embedded across the context, and each test turn requires retrieving the exact multi-hop chain, i.e., all variables assigned to a specific value.
@@
Beyond string retrieval, many real-world long-context applications require semantic understanding, such as retrieving a function from textual descriptions or answering questions from long documents. These tasks are essential in SCBench, as lossy long-context methods often struggle to abstract or comprehend information in multi-request scenarios.
(i) Code.RepoQA: This task requires the model to retrieve a specific function (including its name, input parameters, and full implementation) from a long source code chunk based on a precise natural language description. Unlike the original RepoQA benchmark (Liu et al., 2024c), our inputs extend to 64K tokens, with target functions evenly distributed by position within the codebase. The function descriptions were generated using GPT-4 based on the functions themselves. We also expanded the range of repositories and programming languages to include Python, C++, Java, PHP, Rust, Go, and TypeScript. Each test session involves a GitHub repository, with the model required to retrieve one function per turn across 5 turns.
(ii) En.QA, Zh.QA, En.MultiChoice: These tasks are extended from InfiniteBench (Zhang et al., 2024a), which provides high-quality, human-annotated QA tests based on fictional novels to eliminate external knowledge influence. The tasks require models to locate and process information from lengthy inputs, performing reasoning through aggregation or filtering. The two primary question types are: 1) Aggregation, compiling scattered information across the input (e.g., “How much money in total did A spend on food?”); 2) Filtering, identifying specific details from a larger set (e.g., “What color dress did A wear when A met B for the second time?”). In SCBench, QA pairs sharing the same input context are combined to create shared-context test sessions.
@@
In addition to retrieval, some long-context tasks require leveraging and aggregating global context information, such as summarization, statistical tasks, and in-context learning (ICL) (Yu et al., 2020; Srivastava et al., 2023; Hao et al., 2022). Our benchmark includes three tasks to evaluate how well different long-context methods handle global information in multi-request.
(i) Many-shot ICL: We use datasets from Big-Bench Hard (Srivastava et al., 2023) to evaluate many-shot ICL capabilities. This includes three sub-tasks: date understanding, salient error translation detection, and tracking seven shuffled objects. Many-shot ICL contexts are shared across turns within a test session, and all sub-tasks are presented as multiple-choice questions with four options.
(ii) Math.Find: We extended the math find task from InfiniteBench (Zhang et al., 2024a), expanding it from finding only the maximum value to multiple statistical values. Given a large array, LLMs must find the minimum or median values, requiring effective comprehension of global context, comparisons, and statistical operations.
(iii) En.Sum: This task uses concatenated academic papers from arXiv as input, with document lengths ranging from 8K to 20K tokens. Ground truth summaries, averaging 654 tokens, were generated using GPT-4, which was prompted to produce concise one-sentence summaries for each document. The target documents for each turn are evenly distributed across the full context length.
@@
In real-world applications, LLMs often handle multiple tasks within a single session using a shared input context. For example, users may request both summarization and content retrieval simultaneously. To reflect this, SCBench includes two multi-tasking tasks:
(i) Mix.Sum+NIAH: This task combines document summarization with the Needle in a Haystack (Kamradt, 2023) task using a shared input prompt. A random "needle" is evenly inserted into the input of the En.Sum task (concatenated academic papers). The model alternates between summarization and NIAH retrieval in each test session.
(ii) Mix.RepoQA+KV: This task combines the RepoQA task with KV retrieval using a shared input prompt. Multiple KV pairs are evenly inserted into the RepoQA input (a long chunk of source code), including 100 KV pairs with four target KVs and the rest as distractors. The model alternates between RepoQA and KV retrieval in each test session.